2. Produce truth tables for given Boolean expressions.

i. A̅ B̅ C+A B̅ C̅ + ABC +A̅ B C̅
ii. (A+B̅+C)(A+B+C)(A̅+B+C̅)
1
Expert's answer
2020-01-23T09:46:53-0500

i)

"W= \overline{A}\;\overline{B}\;C, \;
X= A\; \overline{B} \; \overline{C}, \; 
Y= A\;B\;C , \; 
Z= \overline{A}\; B\; \overline{C}"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c:c}
 A & B & C & W& X& Y & 
Z & W+X+Y+Z\\ \hline
 0 & 0 & 0 & 0 &0&0&0&0\\ \hdashline
0 & 0 & 1 & 1 &0&0&0&1\\ \hdashline
0 & 1 & 0 & 0 &0&0&1&1\\ \hdashline
0 & 1 & 1 & 0 &0&0&0&0\\ \hdashline
1 & 0 & 0 &0 &1&0&0&1\\ \hdashline
1 & 0 & 1&0 &0&0&0&0\\ \hdashline
1 & 1 & 0 & 0 &0&0&0&0\\ \hdashline
 1 & 1 & 1 & 0 &0&1&0&1\\
\end{array}"


ii)

"X=A+\overline{B} +C, \;
Y=A+B+C,\;
Z=\overline{A}+B+\overline{C}"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 A & B & C & X& Y & 
Z &XYZ\\ \hline
 0 & 0 & 0 & 1 &0&1&0 \\ \hdashline
0 & 0 & 1 & 1 &1&1&1\\ \hdashline
0 & 1 & 0 & 0 &1&1&0 \\ \hdashline
0 & 1 & 1 & 1 &1&1&1\\ \hdashline
1 & 0 & 0 &1 &1&1&1\\ \hdashline
1 & 0 & 1&1 &1&0&0\\ \hdashline
1 & 1 & 0 & 1 &1&1&1\\ \hdashline
 1 & 1 & 1 & 1 &1&1&1\\
\end{array}"